OpenStudy (anonymous):

CAN ANYBODY SOLVE THIS QUESTION ?? NEED HELP!!! The number of irrational solutions of the equation \[\sqrt{x ^{2}+\sqrt{x ^{2}+11}}+\sqrt{x ^{2}-\sqrt{ x ^{2} +11}} = 4\] is a.) 0 b.)2 c.)4 d.)infinite justify your answer with a proof. (topic - quadratic equations)

OpenStudy (rational):

@dan815

OpenStudy (rational):

@IrishBoy123

OpenStudy (anonymous):

use substitution

OpenStudy (rational):

\[\sqrt{x ^{2}+\sqrt{x ^{2}+11}}+\sqrt{x ^{2}-\sqrt{ x ^{2} +11}} = 4\] Alright! substituting \(\sqrt{x^2+11}=u~~\implies~~ x^2=u^2-11\) the equation then becomes \[\sqrt{u^2+u-11}+\sqrt{u^2-u-11} = 4\]

OpenStudy (anonymous):

you still cant square both sides

OpenStudy (anonymous):

complete the square?

OpenStudy (rational):

im actually thinking of rationalising the left hand side expression

OpenStudy (rational):

but idk how useful it is..

OpenStudy (rational):

completing the square doesn't look that primising..

OpenStudy (anonymous):

you can try

OpenStudy (dan815):

YOU'RE RATIONAL leave to me im irrational

OpenStudy (anonymous):

...

OpenStudy (rational):

you're more like complex/imaginary danny

OpenStudy (rational):

i gave up rationalising half way as it got really complicated

OpenStudy (anonymous):

what do i do now, haha. i dont even go to school.

OpenStudy (anonymous):

we are supposed to solve these questions in less than 2 min -_-

OpenStudy (anonymous):

so there ought to be some trick

OpenStudy (dan815):

ok i just expanded it

OpenStudy (dan815):

is that easier to solve?

OpenStudy (dan815):